5. An auditor recorded the birthdays of 389 Engineering students. Below is the results of the recording (by month). An Engineering student is selected at random (8 marks) MonthChem. Mech.Elect. Other Determine the following probabilities: January February March April May June Jul August September 10 October November December 1 a) the student is born in February 4 17 16 6b) the student is an Electrical Engineer student 16 10 5 c) the student is a Mechanical Engineer student born in August 4 d) the student is either born in July or a Chemical Engineer 0e) Ifthe student was born in October, it is under the Other 5 column 17 6 13 17 18 f) If the student is a Chemical Engineer student, it is not born in March g) Is a Mechanical Engineer student given that the student was born before May h) Is not born in February given that the student is not a Electrical Engineering student. 10 6 0 15 18 18 16 0
Answer #1

a)

total feb students = 8+17+16+6

so proabbaility = (8+17+16+6)/389 = 0.12

b)
total elec eng students (8+16+10+3+4+6+17+7+10+6+0+2)/389 = 0.228

c)

mech eng in aug = 5

so probability is 5/389


d)

born in july = (17+6+17+1)/389 = 0.105

chem eng = (5+8+7+11+4+13+17+18+10+15+6+8)/389 = 0.313

july and chem eng = 17/389 = 0.043

so the

required prob is = 0.105 +0.313 -0.043 = 0.375


please note that we can answer only 4 subparts of a quesiton at a time , as per the answering guidelines